OpenStudy (anonymous):

Find all points (x,y) on graph y=x/x-6 with tangent lines perpendicular to y=6x-5

OpenStudy (xishem):

Try taking the derivative of the function given and work from there.

OpenStudy (anonymous):

I did, I got -6/(x-6) ^2

OpenStudy (xishem):

And what is the slope of the line function that's given? And what slope would a line perpendicular to it have?

OpenStudy (anonymous):

only perpendicular to y=6x-5

OpenStudy (xishem):

And the derivative function represents what?

OpenStudy (anonymous):

what do you mean?

OpenStudy (xishem):

Given a function f(x)...\[f(x)=\frac{x}{x-6}\]what does this represent about the original function, f(x)?...\[\frac{dy}{dx}f(x)=\frac{-6}{(x-6)^2}\]

OpenStudy (anonymous):

the change of the fxn when input change after I got that, I set up the equation -6 / (x-6) ^2 = 6 as the slope of the perpendicular line

OpenStudy (xishem):

Close. 6 would be the slope of a line parallel to y=6x-5.

OpenStudy (anonymous):

then how do I get the actual slope for perpendicular?

OpenStudy (xishem):

The negative reciprocal of one slope is perpendicular to another.

OpenStudy (anonymous):

I got it, solved !!! THanks :)

OpenStudy (anonymous):

can you help me on some other problems?